find the general solution for x(n)and y(n)for x(0)=1 and y(0)=0,
x(n+1)=x(n)+2y(n)
x(n+1)=3x(n)+2y(n) n=0,1,2,3,...


Answer by rapaljer(4671)

find the general solution for x(n)and y(n)for x(0)=1 and y(0)=0,
x(n+1)=x(n)+2y(n)
x(n+1)=3x(n)+2y(n) n=0,1,2,3,...

If n=0, then substitute this into the equation for n:
x(0+1)=x(0)+2y(0)

Since you know the values of x(0) and y(0), substitute these in below:
x(1)= 1 + 2*0
x(1)=1

Now, let n=1, and you have
x(1+1) = x(1) + 2y(1)
x(2) = 1 + 2y(1)

Having said all of this, are you sure you stated this correctly? I see two different formulas for x(n+1), and I don't see any formula for y(n+1). Maybe the problem is not stated correctly. Maybe I just don't know where to go from here.

Dr. Robert J. Rapalje, Retired
Seminole State College of Florida
Altamonte Springs Campus